OpenStudy (anonymous):

The perimeter of a rectangle is 42, and its diagonal is 15. Find its length and width. length (larger side) _____ width (smaller side) _____

OpenStudy (anonymous):

2a+2b=42 a^2+b^2=15^2

OpenStudy (anonymous):

a^2+b^2=225 2(a+b)=42 (a+b)=21

OpenStudy (anonymous):

b=21-a plug that into a^2+b^2=225 a^2+(21-a)^2=225

OpenStudy (anonymous):

The diagonal makes the rectangle into two right triangles. The hypo for each triangle is 15. Lets say that the short side is a and the long side is b. Perimeter= a+a+b+b=42 Pythagorean= a^2 + b^2= 15^2 Simplify perimeter 1st 2a+2b=42 a+b=21 Then simplify the Pythagorean one a^2 + b^2=225 It's easiest to use substitution in this one: make the first one in terms of b (or a), then plug that equation in wherever you see a b in the second equation. a+b =21 b= 21-a a^2+ b^2 = 225 a^2+ (21-a)^2=225 Then solve for a. a^2+ (21-a)(21-a)=225 a^2+ 441 -21a -21a +a^2 =225 2a^2+ 441 -42a =225 2a^2-42a+216=0 Factor that, and when you find out what a is, plug the value into its rightful place in one of the original equations and solve for b.